### v1

# That’s So FETCH: Fashioning Ensemble Techniques for LLM Classification in Civil Legal Intake and Referral

Abstract
Each year millions of people seek help for their legal problems by calling a legal aid program hotline, walking into a legal aid office, or using a lawyer referral service. The first step to match them to the right help is to identify the legal problem the applicant is experiencing. Misdirection has consequences. Applicants may miss a deadline, experience physical abuse, lose housing or lose custody of children while waiting to connect to the right legal help. We introduce and evaluate the FETCH classifier for legal issue classification and describe two methods for improving accuracy: a hybrid LLM/ML ensemble classification method, and the automatic generation of follow-up questions to enrich the initial problem narrative. We employ a novel data set of 419 real-world queries to a nonprofit lawyer referral service. Ultimately, we show classification accuracy (hits@2) of 97.37% using a mix of inexpensive models, exceeding the performance of the current state-of-the-art GPT-5 model. Our approach shows promise in significantly reducing the cost of guiding users of the legal system to the right resource for their problem while achieving high accuracy.
keywords:
[…]
Introduction
Finding the right lawyer for your legal problem can be hard. Those who know a lawyer may discover that they practice a different kind of law than they need. Legal-aid eligible populations may spend significant time identifying the right program to apply to. Others, over the income limits for legal aid, may spend significant time asking friends or call a lawyer referral service.
Getting legal triage or referral wrong means that the applicant may be prevented from getting legal help that they are otherwise entitled to, with consequences as severe as loss of liberty, inability to get protection from domestic violence, loss of housing, or loss of custody of a child; or as simple as loss of time and money speaking to an attorney who cannot help them with their specific legal problem.
One tool for classifying natural language text is a large language model (LLM). Even current state-of-the-art LLMs do not classify with 100% accuracy. “Ensemble” classification, which combines multiple classification methods to produce a single combined score [7], is a method of improving classification accuracy.
Ensemble classification methods may use multiple models in series, with routing (where a less powerful model routes the text to the model that performs best on that kind of text) [10], or using a voting method where all models generate a score that is combined using an algorithm [9]. This paper introduces the FETCH (Fast Ensemble Tagging and Classification Helper) classifier. FETCH is a multi-label weighted ensemble voting classifier that combines LLMs, keyword matching, and traditional ML.
This work was produced in collaboration with the Virginia Legal Aid Society and the Oregon State Bar. Virginia Legal Society (VLAS) is a nonprofit legal aid program that serves the southern portion of the U.S. state of Virginia and receives approximately 18,000 calls to human intake workers each year. The Oregon State Bar (OSB) is the nonprofit mandatory state-wide bar for the US State of Oregon and fields approximately 100,000 inquiries to its bar referral service each year. For several years, they have used a simple online webform to gather information from applicants. Trained staff then manually match applicants to the appropriate legal referral.
Ultimately, we show that a combination of 3 inexpensive LLM models: GPT-5-nano, gemini-2.5-flash, and mistral-small, together with a simple keyword matching system and the traditional ML Spot API [8] can exceed the performance of the current state-of-the-art GPT-5 model at a significant cost and speed savings.
0.1 Research questions
  • RQ1: Can small-model ensembles deliver equivalent or superior classification accuracy to frontier models while reducing cost in legal intake?
  • RQ2: What types of legal intake cases remain most challenging for ensemble and LLM classifiers, and how can error analysis inform classifier design?
1 Prior work
1.1 AI for access to justice
The application of AI to the domain of access to justice is well established and is now the topic of a biannual workshop co-located with JURIX and ICAIL [15], as discussed extensively in [21]. See, e.g., applications to legal information tools [6, 26], form-filling and expert systems [20, 19, 25], and dispute resolution [24, 4, 5].
1.2 Ensemble classification of natural language text
Ensemble classification methods, where multiple models are used in sequence or at the same time, have been used for some time to improve accuracy in classification prior to the widespread popularity of large language models, popularized in [7] . See, e.g., [12], discussing early neural network and decision tree classification. LLMs have been applied to ensemble classification in diverse domains, including psychology [14] and law [10, 13, 17].
1.3 Legal intake and advice
[21] discussed the application of LLMs to the task of legal intake, focusing specifically on threshold eligibility determinations in the legal aid context. The work in this paper builds on and generalizes that prior work to the broader legal classification task.
2 The legal problem classification task
An attempted legal referral can end in one of three ways:
  • The applicant is matched with a provider who specializes in the kind of problem the applicant is experiencing.
  • The intake worker may decide the applicant has a problem that does not have a legal solution, due to inapplicability, statutes of limitation, or jurisdiction.
  • The intake worker may decide that the applicant is not experiencing a legal problem at all, but instead may be under a delusion. [^1] A classic, commonly experienced example described by staff at the Oregon State Bar was “I have a microchip in my head”. The phenomenon of body-control delusions has a well-documented link to paranoid schizophrenia dating to 1919. See, e.g., [22]. In these cases, the intake worker may pay extra attention to handle the applicant’s problem with sensitivity.
Classification is central to each of these tasks. And leaving the classification up to the applicant is risky. Applicants may have an idea of the kind of lawyer they need to help with their problem, but that idea may be wrong. For example, a layperson might be shown all 244 categories and then be asked to decide that their problem getting Social Security Disability benefits falls under the category “Administrative Law” (a relatively obscure term for a layperson). Informally, we observed applicants struggle when faced with distinguishing between abstract legal categories, and this is supported by studies such as [27] (showing that only 29% of people experiencing a legal problem recognize it as such) and [11] (emphasizing that self-assessed legal understanding often diverges from actual understanding).
Perfect referral is not always possible in a given geographical area, and applicants may settle for a next-best match in those cases.
3 The FETCH classifier
The FETCH classifier is a Python FastAPI application available as a REST-like API. The API is currently consumed by two separate applications: the Oregon State Bar’s online lawyer referral service, and the Virginia Legal Aid Society’s phone-based intake system.
We employ the same prompt with 3 LLMs, which are each prompted with the taxonomy and the applicant’s problem description to determine if:
  1. 1.
    The problem can be classified into one or more of 244 taxonomy nodes.
  2. 2.
    More information is needed to accurately classify the applicant’s problem.
  3. 3.
    The applicant has not described a legal problem at all.
We experimented with different levels of detail in the taxonomy for the prompt, but settled on an abbreviated level, with simply the name of each category and its parent category.
The results are then scored by the number of times each node appears in the results of one of the models. The models are manually weighted according to past performance and results are ranked accordingly. Finally, the top 2 nodes are shown to the applicant.
If the ensemble does not have sufficient confidence in the classification, each LLM model in the ensemble is instructed to generate 3 follow-up question candidates. The top 3 questions (after semantic merging with the assistance of GPT-5-nano) are then shown to the applicant.
4 Oregon State Bar’s referral form
The first implementation of the FETCH classifier is in the Oregon State Bar’s online referral system (Figure 1). Applicants enter the information about their legal problem in a standard web form. Follow-up questions continue to use a form metaphor rather than simulated chat. The LLM is instructed to provide questions with radio buttons and checkboxes as much as possible, which are rendered by the frontend application. We hypothesize that the form both reduces effort for the applicants and reduces the “uncanny valley” [23] effect that users experience when talking to an LLM [16].
Figure 1: Screenshot of the lawyer referral service with integrated FETCH classifier, showing an automatically generated follow-up information form. Here, the applicant has entered an ambiguous query: “I am getting kicked out” and the FETCH classifier generates 3 questions to clarify if this is an eviction or other civil matter.
Because it generates follow-up questions automatically, the FETCH classifier enables the tool to gather more information without a human having to call the applicant.
Applicants who are screened out are given the option to call a human intake worker at the Oregon State Bar.
5 Experimental design
5.1 Dataset
In our experiment, we employed a novel dataset of 419 human-annotated queries made to the Oregon State Bar Referral Service over a period of a few weeks in 2025. The dataset contains a total of 31,716 words, with a mean problem description length of 74 words. Approximately 300 of the queries were annotated to one of the 15 top levels of the legal taxonomy only, while the 119 remaining queries were annotated to a specific “terminal node” in the OSB taxonomy. Annotation was performed by legal professionals at the Oregon State Bar during the regular operation of their referral program. The dataset was anonymized by the bar referral service prior to our analysis.
Figure 2: Distribution of problem type by top-level category name in the dataset.
Queries in the data set show a diversity in length (Fig. 3) and detail. Spelling errors and idiosyncratic punctuation appear throughout.
Some examples from the dataset:
  • Need bankruptcy lawyer
  • Major car accident - DUII
  • I Have a small sole proprietor owned business in construction (tubs&counterrs) & cleaning. Im having issues being paid. Im looking to find a lawyer that knows how to… do contracts and lien property filing for my company. Im just starting to get a profit for the first time. Now I can’t seem to be paid. Invoices over due. (Approx 50% of them) causing me to do this….. if i could get a Refural please.
    High-stakes problems, from stalking and domestic violence abuse to emergency housing, are well represented in the dataset: [^2] Unlike in [21], we did not observe any instances of model “censorship” or refusal to classify situations involving violence or abuse, a promising indication of the suitability of LLMs for problems in the legal domain.
  • Hello, Here is the information, I have been served with an eviction notice with numerous months of non payment, for several reason in which i won’t go into detail at this time , until i have met with a lawyer.Iam currently unable to work.It is very important that i find a reputable legal aid lawyer. This is an urgent request in which in hope it will be resolved soon.I have a family emergency which is a life threatening illness to attend to […]
  • I was granted a temporary stalking order on Friday, [date]. A contested hearing is scheduled for [date] and I need an attorney to represent me as a petitioner. I would also like further legal advice on taking legal action against someone who is stalking me. The respondent to my stalking order initially made contact with as my client in a job I was working as social services provider. Months after I had stopped working with her as a housing advocate, she […]
Figure 3: Distribution of applicant query lengths in the dataset.
5.2 Taxonomy
Our study uses the 244 node taxonomy built by the Oregon State Bar. The Oregon State Bar operates a general-purpose bar referral service covering all civil and criminal case types. The Oregon State Bar’s taxonomy is hierarchical with two levels, with the top level grouping the terminal nodes into 15 categories, such as “Administrative Law”, “Real Property”, and “Criminal Law”. Terminal nodes include more specific classification, such as “Family Law > QDROs (Qualified Domestic Relations Orders)”, “Real Property > LLTN – ADA issues/Animals/Accommodations”, “Bankruptcy > Farms”. Importantly, the topics covered by the taxonomy are truly national in scope, representing a superset of alternative taxonomies such as the 135-node Legal Matter Specification Standard (commonly referred to as SALI taxonomy) [18]. We also used the results from the Spot API [8], which use the conceptually smaller civil legal aid-focused LIST Taxonomy. [^3] https://taxonomy.legal/
5.3 Classifiers
Our study encompasses 9 separate classification approaches:
  • A simple keyword-based classifier.
  • A term frequency/inverse document frequency classifier.
  • GPT-4.1-mini and GPT-4.1-nano.
  • Gemini-2.5-flash.
  • Mistral-small-3.2-24b-instruct.
  • Suffolk Legal Innovation and Technology Lab’s traditional ML Spot [8] classifier (build 10).
  • GPT-5 and GPT-5-nano, released near the end of our study.
The ensemble method was a simple weighted vote, with weights calculated based on the raw performance of each classifier across the training dataset. The classifier was limited to providing no more than the 2 “best” labels based on the weighted ranking.
We are able to use the additional signal of a confidence level from the SPOT classifier. For Gemini models, we use logprobs as returned in the LLM’s response as a proxy for confidence level, as suggested by [10]. [^4] Unfortunately, the GPT-5 model series from OpenAI no longer makes logprobs available.
5.4 Evaluation method
We evaluated the raw performance of the classifier with the help of the open-source Promptfoo tool against the human-annotated dataset with a hits@2 scoring system. We scored a result as “pass” if the human-annotated label was present in the response, and scored it as passing with a score of 0.5 if at least the top-level node was present in the labels.
6 Prompt iteration
At the initial development stage for the prompt, we experimented with different levels of detail in the prompt, settling on simply including the taxonomy labels without additional explanation.
7 Results
7.1 Hits@2 results
Table 1 shows the results as scored using hits@2. Notably, the results show that the ensemble method performs best, exceeding the performance of best-of-class GPT-5 alone and greatly exceeding the average performance of each inexpensive model used in the ensemble classifier. GPT-5-nano’s strong performance at a fraction of the cost of GPT-4.1-mini means that we pivoted to including it in the ensemble classification shortly after its release.
Method Hits@K (top 2 labels) (%)
Spot [8] 59.43
Keyword 54.18
TF-IDF 31.03
Gemini 2.5-flash 87.35
Mistral small 87.83
GPT-4.1-mini 81.38
GPT-5-nano 87.11
GPT-5 96.66
Ensemble (GPT-5-nano, Gemini 2.5-flash, Mistral small, Keyword, Spot) 97.37
Table 1: Model and ensemble accuracies.
7.2 Cost efficiency
An important justification for the use of weighted ensemble voting is cost efficiency. As shown in Table 2, the ensemble voting method outperforms the state-of-the-art GPT-5 model on a cost basis by a factor of more than 3.
Model Cached Uncached Output Avg. cost Est. annual
(\$/1k) (\$/1k) (\$/1k) (\$/1k req) cost 100k (\$)
mistral-small-3.2-24b-instruct 0.026 0.008 0.030 0.064 6.40
gemini-2.5-flash 0.039 0.048 0.750 0.837 83.75
GPT-5 0.065 0.201 3.002 3.267 326.71
GPT-5-nano 0.003 0.008 0.120 0.131 13.07
GPT-4.1-nano 0.013 0.016 0.120 0.149 14.91
GPT-4.1-mini 0.052 0.064 0.480 0.596 59.64
Table 2: Cost model assumes a typical query of 519.5 cached input, 160.5 uncached input, and 300.2 output tokens. Provider assumptions: OpenAI cached input is priced at 10% of the input rate; Gemini 2.5 Flash cached input is \$0.075 per 1M tokens; OpenRouter mistral-small-3.2-24b-instruct treats cached input at the same price as input. Note that the Spot API is free for nonprofits.
The latency of the small LLMs is also much lower. For example, GPT-5 has a minimum latency of over 5 seconds [1] compared to GPT-5-nano’s minimum latency of 2.2s [2].
7.3 Errors
We closely examined the 11 errors of the ensemble classification method. As in [21], we discovered probable human annotation errors through the LLM classification, totaling four such errors. We also discovered two duplicate entries (omitted in Table 1) in the human annotated dataset which may overweight the LLM classification errors. [^5] We leave these annotation errors in our tops@2 score for consistency. The full error results are summarized in Table 1.
Short summary Human annotation LLM Error classification
Roofing replacement contract; contractor obligations/next steps Real Property > Construction/Contractors General Litigation , Consumer Law LLM error
Neighbor harassment/stalking (incl. bear spray); safety concerns Real Property > Neighbor Disputes/Nuisance General Litigation > Neighbor Disputes/Nuisance, General Litigation > Personal Injury Synonymous
Dispute over dog ownership/possession with ex-partner; access withheld Consumer Law General Litigation > Animal Law, Family Law > General (Divorce/Separation) Ambiguous
Wedding planner using event photos/video for promotion; cease-and-desist requested Consumer Law General Litigation > General Torts/Privacy, Intellectual Property > Trademark/Copyright Ambiguous
Lost trial for mobile-device while driving; seeks appeal Consumer Law Criminal Law > Post Conviction/ Appeals, Criminal Law > Traffic Offenses Human annotation error
Wants help filing a small-claims case Criminal Law Consumer Law > Small Claims Advice, General Litigation Human annotation error
Car accident; likely uninsured; facing costs; asks about payment plans Debtor/Creditor General Litigation > Personal Injury, General Litigation > Property Damage LLM error
Pandemic unemployment benefits issue; denial/overpayment and appeal options Debtor/Creditor Administrative Law > Unemployment, Labor & Employment > General Human annotation error
Physician Assistant practice without supervising MD/DO; licensing/supervision compliance Labor & Employment Administrative Law > Professional Licensing, Business & Corporate > General (Contracts, Business, Organization) Human annotation error
Table 3: Detailed error results.
For the two errors we classified as clear “LLM error,” the LLM response was a form of “general litigation” instead of a more specific category. FETCH can optionally generate follow-up questions, and in four of the error results it did generate apparently relevant questions.
8 Discussion
8.1 RQ1: Can small-model ensembles deliver equivalent or superior classification accuracy to frontier models while reducing cost in legal intake?
We have shown that small-model ensembles substantially improve on the performance of older machine learning and keyword-matching approaches to classifying legal problems, and slightly improve on the performance of frontier LLM models. What’s more, they do so with lower latency and at a significantly lower cost, at least $\tfrac{1}{3}$ of the cost of using a frontier model for the same task.
The LLM-enabled classification approach shows promise to offer nonprofit bar referral services and legal aid programs a cost-effective way to deliver high quality referral or eligibility determination information to applicants for legal help.
8.2 RQ2: What types of legal intake cases remain most challenging for ensemble and LLM classifiers, and how can error analysis inform classifier design?
We did not observe any patterns in the errors in our ensemble classifier’s results. They did not appear to be biased towards one type of legal problem over another. We did not observe a strong pattern of the LLM failing to distinguish two similar categories. Importantly, this tells us that enriching the category descriptions or adding additional clarifying instructions in the prompt is not likely to significantly improve classification performance.
The number of apparent human annotation errors together with the ambiguous classifications suggest that the baseline human performance at the classification task may be comparable to the LLM ensemble classification method.
Although we did not directly measure the impact of asking follow-up questions, we theorize that it would further improve LLM classification accuracy.
Overall, the errors made by the LLM appear to be remarkably few (approximately 2%), and consistent with the kind of challenges humans may face in similar circumstances.
8.3 Limitations and risks
The results of the study are strong. However, there are some limitations:
  • We did not extensively test the “no legal problem” condition, particularly when the applicant query shows signs of delusion. We have thoughts on how to sensitively escalate such cases to a human that require further testing.
  • We did not examine the real world benefit to an applicant to having quick and accurate classification of their legal problem and immediate referral without the delay of human classification. For example, if this led to faster resolution of the applicant’s problem.
  • Whenever an LLM is involved in questions of judgment, embedded biases [3] must be considered. We note that the strong agreement between human and LLM in this experiment, as well as the diversity of methods and models in the ensemble, reduces this concern. In addition, in the systems we deployed, applicants always have an opportunity for human review.
9 Conclusion and future work
The ensemble method shows promise to improve performance on a high-stakes classification task. Importantly, our results show that a weighted combination of inexpensive small language models can outperform the current state-of-the-art model, GPT-5, at approximately $\tfrac{1}{3}$ of the cost.
Although the error rate is relatively small in our study, future work could further narrow the gap. A routing method based on perceived confidence or disagreement among the model responses could supplement the inexpensive LLM models with a call to a best-of-class LLM such as GPT-5. We did not study the impact of asking follow-up questions in this experiment, but that bears future testing as well. An approach that has human annotators manually remove critical information from answers would be labor intensive but a solid way to answer this question.
Finally, we always look for information about human baseline performance when deciding if an LLM’s performance is “good enough”. We look forward to finding a way to gather this information in a future study.
Acknowledgments
The author thanks Virginia Legal Aid Society and Oregon State Bar for funding and data access. Additionally special thanks to David Colarusso, Jack Adamson, and Jim Graszer for their contributions and support throughout this project.